OpenStudy (anonymous):

Someone want to complete the square for me, on the bottom? Never truly learned how to do it... know you have to take half the mid. term.... Integral of ((1/(9x^2 - 36x + 40)

hartnn (hartnn):

what is middle term ?

OpenStudy (anonymous):

I lied.... now I think it's supposed to change to 9(x -2)^2 + 4..... which doesn't help me anyway.... because I'm not sure how to get that into the 1/(x^2 +1)....

hartnn (hartnn):

take 9 common from 9(x -2)^2 + 4 what u get ?

OpenStudy (anonymous):

Well.... techincally I'm supposed to get the four out of there first. so it's 1/4 times the integral.... and based on notes it's almost like I would have to go 9/4(x-2)^2 + 1...... which is just.... ugh.

OpenStudy (anonymous):

I don't think I took good notes on this junk.

hartnn (hartnn):

ok, u can do that also. so u have 9/4(x-2)^2 +1 = \(\huge ((2/3)(x-2))^2+1 \) now put y= 2/3(x-2) and tell me what u get

hartnn (hartnn):

from 9/4 (2/3)^2 = 9/4 i wanted to bring 9/4 inside square term

hartnn (hartnn):

confused ?

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

I'm still confused on how the (2/3)^2 is equivalent to 9/4..... because.... it's not..... or else I'm just worse at math than I thought. I would have reversed the two numbers...

hartnn (hartnn):

omg! i id a typo, sorry, its (3/2)^2 sorry.

hartnn (hartnn):

one more typo *did

OpenStudy (anonymous):

Okay, I think I can get the right answer from there, thanks! .... but I'll probably be right back on the site with more integral crap.

hartnn (hartnn):

sure do come, happy to help ^_^